# Weakly supervised anomaly detection with event-level variables

### v1

@@ IV Methods
We select the two highest $p_{T}$ $\tau$-tagged jets with |$\absolutevalue{\eta}<2.4$ as our resonance candidate. The leading $\tau$ jet is required to have $p_{T}>40$ GeV. We use the $m_{\mathrm{vis}}$ as our resonance variable in the CATHODE approach and require $m_{\mathrm{vis}}>100$ GeV. For simplicity, we consider a single signal region (SR) with a mass window of 175–300 GeV, which contains the majority of our generated signals, as visible in Fig. 1. In a real search, a scan could be performed repeating the procedure for different signal window choices to cover the full range as has been done in previous anomaly detection searches.
@@ IV.1 Features
To establish a consistent representation of the phase space manifold, we require fixed dimensionality of objects across all events. We define events in terms of constituent objects (jets, electrons, muons, or photons), with each event represented by exactly fifteen objects. For events with fewer than fifteen physical objects, we perform zero-padding with four-vectors as $(E,p_{x},p_{y},p_{z})=(0.1,0.1,0,0)$. Objects are ordered by decreasing transverse momentum ($p_{T}=\sqrt{p_{x}^{2}+p_{y}^{2}}$) to ensure consistent event representation.
The fifteen objects yield fifteen simplex coordinates per event, following the manifold construction described in Section II. Our analysis reveals that the first ten simplex coordinates capture the majority of the relevant phase space topology, with the remaining coordinates primarily corresponding to the zero-padded objects in typical events.
To complement the phase space variables, we incorporate event-level kinematic observables: the scalar sum of non-$\tau$ jet $p_{T}$, the missing transverse momentum (MET), and the angular separation ($\Delta$R $=\sqrt{(\Delta\eta)^{2}+(\Delta\phi)^{2}}$, where $\eta$ is the pseudorapidity and $\phi$ is the azimuthal angle) between the two $\tau$ jets. These features encode global event characteristics that are not fully captured by the phase space coordinates alone. Our final feature vector consists of: the first ten simplex coordinates, MET, non-$\tau$ jet $p_{T}$ sum, and di-$\tau$ $\Delta$R.
Because of the unique characteristics of the phase space coordinates, it was found to be necessary to add a small amount of noise to the zero-padded phase space coordinates so that they could be well simulated by the generative model. Details of this preprocessing are given in Appendix A.1.
@@ IV.2 Classification and Background Evaluation
Our generative model uses a normalizing flow architecture based on rational quadratic splines [76], and is trained in the sidebands. For simplicity, the training is performed on a background-only sideband sample, assuming the signal contamination to be negligible. A kernel density estimate is used to learn the $m_{\mathrm{vis}}$ distribution for use in synthetic background generation. Details of the model architecture and exact generation procedure are given in Appendix A.2.
@@
For the weakly supervised classification task, we employ an ensemble of 50 histogram-based gradient boosted decision trees (BDTs). Ensembles of BDTs have demonstrated superior performance over neural networks in weakly supervised training with high-dimensional feature spaces [46]. Our implementation utilizes the HistGradientBoosting classifier from scikit-learn [77].
@@ IV.3 Synthetic Background Quality Validation
A comparison of the synthetic events generated by the background model and the true background events in the SR is shown in Fig. 3.
Figure 3: A comparison of the distributions of true background events in the signal region and the synthetic ones generated by our normalizing flow. Simplex coordinate 10 is omitted for brevity, but has similar distributions.The last bin in each histogram shows overflow entries beyond the range of the histogram axis. The flow is seen to model most of the distributions well. Some discrepancies are seen in the modeling close to variable boundaries at zero.
The synthetic events mimic the background relatively well, besides some residual discrepancies seen for the lower values of each distribution. Normalizing flows are known to have difficulties modeling sharp features near the edges of distributions. As this region of phase space is the most background-dominated, its mismodeling may have a reduced impact on signal identification performance.
To validate the quality of our synthetic background, we applied our BDT classifier to distinguish real background events from synthetic ones. This process yields a classification area under curve of 0.55, close to the random-chance value of 0.5. Imperfect modeling of the background may dampen sensitivity in the CATHODE approach, as the weakly-supervised classifier may focus on the differences between the synthetic and true backgrounds rather than identifying signals. However, even with our imperfect background modeling, we have seen strong anomaly detection performance with our chosen signals. We therefore conclude our background model is of sufficiently high quality for the anomaly detection procedure to work, and leave further improvements to future studies.
@@ V Results
Figure 5: A summary of the sensitivity enhancement of our anomaly detection approach. The x-axis shows the size of the signal, quantified as $\frac{S}{\sqrt{B}}$ prior to any anomaly detection selection. The y-axis shows the approximate signal significance achieved after the application of an anomaly score cut that is 1% efficient on the background. Our procedure is able to enhance the signal significance above the discovery threshold ($5\sigma$) for moderately sized signal injections $\leq 2\sigma$ for all considered signal models.
@@ VI Conclusion
We have demonstrated the potential for weakly supervised anomaly detection to be applied to a new class of resonance searches. Di-object plus X searches, in which one looks for a resonance decaying to two standard model particles, but produced in association with other anomalous event activity, represents a new class of searches ripe for the application of anomaly detection. In this exploratory work, we focused on a di-$\tau$ plus X final state. We have employed new physically-motivated phase space variables to robustly capture salient features of multi-particle kinematics. We demonstrated that our anomaly detection approach can reach discovery-level significances for signals that would be missed in a conventional bump-hunt approach. Further work can be done to improve the generative modeling of topologically interesting variables like the hypersphere phase space coordinates. We could potentially gain sensitivity from exploring other event-level variables that quantify object type to complement the kinematic information of the phase space variables. It may be fruitful to explore non-Euclidean normalizing flows, which may better capture the topological features of the phase-space variables, or diffusion models which have been shown to have state of the art performance at physics-based generative tasks [78].
@@ Acknowledgments
We would like to thank Joe Incandela, Tianji Cai and Nathaniel Craig for useful discussions. We use the SK_CATHODE [79] library for implementation. We used the UCSB computational facilities administered by the Center for Scientific Computing at the California NanoSystems Institute and Materials Research Laboratory (an NSF MRSEC; DMR-1720256) and purchased through NSF CNS-1725797. UCSB is supported by the US Department of Energy under grant DE-SC0011702. Support for UCSB is also made possible by the Joe and Pat Yzurdiaga endowed chair in experimental science. LB was also supported by the National Science Foundation Graduate Research Fellowship Program under Grant No. (NSF 2139319). OA is supported by FermiForward Discovery Group, LLC under Contract No. 89243024CSC000002 with the U.S. Department of Energy, Office of Science, Office of High Energy Physics. LM and MS acknowledge support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC2121 “Quantum Universe” – 390833306 as well as by the Bundesministerium für Bildung und Forschung under project 05H24GUA. SS and PM are supported by NSF Grant PHY2310072.
@@ A.1 Preprocessing
In the CATHODE method, the synthetic background samples generated by the normalizing flow should be indistinguishable from the true background events in the SR. If this is not the case, then the weakly supervised classifier will fail to learn to identify the signal, instead learning to distinguish the real and synthetic background events. The use of phase space variables, which have a distinct topological structure and complex interdependence, made this non-trivial to achieve.
A key challenge in the feature engineering was the handling of zero-padded object coordinates in the phase space. Because of normalization conditions, the phase-space simplex variables of the zero-padded objects have a distinct non-zero value that is exactly the same for all objects in the event, but the value varies between events due to a dependence on the other event objects. This strict non-trivial structure is very difficult for the normalizing flow to model exactly. Without proper preprocessing, we found that a classifier would trivially distinguish between real and synthetic background events based on very small variations in this regularity, undermining the CATHODE methodology.
We implemented a stochastic noise addition procedure that carefully applied a small amount of Gaussian noise (approximately 1% of the coordinate magnitude) to the simplex coordinates of padded objects. This technique effectively blurred these strict regular features of the zero-padded objects, allowing for high-quality modeling by the generative model, without distorting the overall distribution characteristics.
Following this noise addition, we employ a sequence of transformations to optimize the feature space for normalizing flows. First, all features are shifted and scaled to the range $x\in(0,1)$ Then a logit transformation ($\text{logit}(x)=\ln\left(\frac{x}{1-x}\right)$) is applied to convert bounded domains into unbounded ones. The transformed features are then standardized by subtracting the mean and dividing by the standard deviation of the training set, respective to each feature. This preprocessing pipeline ensures that the normalizing flow operates on a well-conditioned feature space while preserving the essential structure of the phase space manifold.
@@
Drell–Yan + $t\bar{t}$ $\sim$158,000 $\sim$127,000 $\sim$31,000
Training Validation Testing
@@ A.2 Generative Model
Our generative model is based on a normalizing flow architecture. It consists of 6 layers of a spline-based masked piecewise rational quadratic autoregressive flow [76]. Each layer uses 6 MADE blocks, a hidden feature size of 512, 24 bins, and a tail bound of 6. The flow is trained with a learning rate of $5\times 10^{-5}$ using the Adam optimizer [80] and a batch size of 256. The model is trained until the validation loss fails to improve for 50 epochs.
The flow is trained on events from the sideband regions, conditioning on the di-$\tau$ visible invariant mass. The flow learns a bijective, invertible mapping $z=f(\vec{x};m_{vis})$ between the feature space $x$ and a latent space $z$ following a simple base distribution (unit normal). The power of this approach is that once trained, the inverse function $\vec{x}=f^{-1}(z;m_{vis})$ can be queried for any mass value $m_{vis}$, including those in the SR.
After training, we select the model state with the lowest validation loss for the subsequent interpolation and sampling steps.
To learn the di-$\tau$ mass distribution in the SR, we utilize a kernel density estimate (KDE) implemented from the scikit-learn library [77]. We use a Gaussian kernel with a bandwidth of 0.01. The KDE is fit to the mass distributions in the sideband training set and then a mass distribution for the SR is generated.
To generate synthetic background events in the SR, we first sample mass values from a kernel density estimate of the SR mass distribution. We then draw samples from the base distribution $z\sim\mathcal{N}(0,1)$ The inverse transformation $\vec{x}=f^{-1}(z;m_{vis})$ learned by our flow is then applied to generate synthetic background events.

### v4

@@ IV Methods
[…]
@@ IV.1 Features
To establish a consistent representation of the phase space manifold, we require fixed dimensionality of objects across all events. We define events in terms of constituent objects (jets, electrons, muons, or photons), with each event represented by exactly fifteen objects. For events with fewer than fifteen physical objects, we perform zero-padding with four-vectors as $(E,p_{x},p_{y},p_{z})=(0.1,0.1,0,0)$. We avoid four-vectors with exactly zero momentum because this leads to division by zero in the phase space coordinate calculation, causing numerical instabilities. Objects are ordered by decreasing transverse momentum to ensure consistent event representation.
The fifteen objects yield fifteen simplex coordinates per event, following the manifold construction described in Section II.
Our analysis reveals that the first ten simplex coordinates, which corresponds to the ten highest $p_{T}$ objects, capture the majority of the relevant phase space topology, with the remaining coordinates primarily corresponding to the zero-padded objects in typical events.
[…]
@@ IV.2 Classification and Background Evaluation
Our generative model uses conditional flow matching [85, 86, 87], and is trained in the sidebands. For simplicity, the training is performed on a background-only sideband sample, assuming the signal contamination to be negligible. A kernel density estimate is used to learn the $m_{\mathrm{vis}}$ or $m_{\mu\mu}$ distribution for use in synthetic background generation. Details of the model architecture and exact generation procedure are given in Appendix A.2.
@@
For the weakly supervised classification task, we employ an ensemble of 50 histogram-based gradient boosted decision trees (BDTs). Ensembles of BDTs have demonstrated superior performance over neural networks in weakly supervised training with high-dimensional feature spaces [55]. Our implementation utilizes the HistGradientBoosting classifier from scikit-learn [88]. Hyperparameters for the BDT are kept to be the defaults as in the SK_CATHODE [89] repository.
@@ IV.3 Synthetic Background Quality Validation
We mix two well-performing generated backgrounds in equal proportions, as each may capture different aspects of the background distribution. A comparison of the synthetic events generated by the background model and the true background events in the SR is shown in Fig. 3.
Figure 3: A comparison of the distributions of true background events in the signal region and the synthetic ones generated by our normalizing flow for di-$\mu$. Simplex coordinate 10 is omitted for brevity, but has similar distributions.The last bin in each histogram shows overflow entries beyond the range of the histogram axis. The flow is seen to model most of the distributions well. Some minor discrepancies are seen in the modeling close to variable boundaries at zero. The feature distributions and agreement between synthetic and true are similar in the di-$\tau$ channel.
To validate the quality of our synthetic background, we applied our BDT classifier to distinguish real background events from synthetic ones in the signal region. The BDT classifier uses same architecture and features as are used in the weakly supervised training. The training corresponds to what the weakly supervised classifier would see if there were only background events in the signal region. A near-random classification performance shows that such a classifier cannot tell the samples apart. This indicates that if there was instead signal events in the signal region the classifier would focus on correctly classifying them to improve the loss, rather than concentrating on the differences between the synthetic and true background. This test yields a classification area under curve of 0.52 for both channels, close to the random-chance value of 0.5.
Imperfect modeling of the background may dampen sensitivity in the CATHODE approach, as the weakly-supervised classifier may focus on the differences between the synthetic and true backgrounds rather than identifying signals. However, even with our imperfect background modeling, we have seen strong anomaly detection performance with our chosen signals. We therefore conclude our background model is of sufficiently high quality for the anomaly detection procedure to work, and leave further improvements to future studies.
@@ V Results
Figure 5: A summary of the sensitivity enhancement of our anomaly detection approach for both di-$\mu$ (left) and di-$\tau$ (right). The x-axis shows the size of the signal, quantified as $\frac{S}{\sqrt{B}}$ prior to any anomaly detection selection. The y-axis shows the approximate signal significance achieved after the application of an anomaly score cut that is 1% efficient on the background. Our procedure is able to enhance the signal significance above the discovery threshold ($5\sigma$) for moderately sized signal injections $\leq 2\sigma$ for all considered signal models.
@@ VI Conclusion
We have demonstrated the potential for weakly supervised anomaly detection to be applied to a new class of resonance searches. Di-object plus X searches, in which one looks for a resonance decaying to two standard model particles, but produced in association with other anomalous event activity, represents a new class of searches ripe for the application of anomaly detection. In this exploratory work, we focused on di-$\tau$ and di-$\mu$ plus X final states. We have employed new physically-motivated phase space variables to robustly capture salient features of multi-particle kinematics. We demonstrated that our anomaly detection approach can reach discovery-level significances for signals that would be missed in a conventional bump-hunt approach. Further work can be done to improve the generative modeling of topologically interesting variables like the hypersphere phase space coordinates. We could potentially gain sensitivity from exploring other event-level variables that quantify object type to complement the kinematic information of the phase space variables. It may be fruitful to explore non-Euclidean normalizing flows, which may better capture the topological features of the phase-space variables.
@@ Acknowledgments
We would like to thank Joe Incandela, Danyi Zhang and Sanjit Masanam for useful discussions. We use the SK_CATHODE [89] library for implementation. We used the UCSB computational facilities administered by the Center for Scientific Computing at the California NanoSystems Institute and Materials Research Laboratory (an NSF MRSEC; DMR-1720256) and purchased through NSF CNS-1725797. UCSB is supported by the US Department of Energy under grant DE-SC0011702. Support for UCSB is also made possible by the Joe and Pat Yzurdiaga endowed chair in experimental science. LB was also supported by the National Science Foundation Graduate Research Fellowship Program under Grant No. (NSF 2139319). OA is supported by FermiForward Discovery Group, LLC under Contract No. 89243024CSC000002 with the U.S. Department of Energy, Office of Science, Office of High Energy Physics. LM and MS acknowledge support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC2121 “Quantum Universe” – 390833306 as well as by the Bundesministerium für Bildung und Forschung under project 05H24GUA. SS and PM are supported by NSF Grant PHY2310072. TC gratefully acknowledges the support of the SCale Family Fund during the career transition period.
@@ A.1 Preprocessing
In the CATHODE method, the synthetic background samples generated by the normalizing flow should be indistinguishable from the true background events in the signal region. If this is not the case, then the weakly supervised classifier will fail to learn to identify the signal, instead learning to distinguish the real and synthetic background events.
We employ a sequence of transformations to optimize the feature space for our generative model. First, all features are shifted and scaled to the range $x\in(0,1)$ Then a logit transformation ($\text{logit}(x)=\ln\left(\frac{x}{1-x}\right)$) is applied to convert bounded domains into unbounded ones. The transformed features are then standardized by subtracting the mean and dividing by the standard deviation of the training set, respective to each feature. This preprocessing pipeline ensures that the normalizing flow operates on a well-conditioned feature space while preserving the essential structure of the phase space manifold.
@@
$\tau\bar{\tau}$ Drell–Yan + $t\bar{t}$ $\sim$158,000 $\sim$127,000 $\sim$31,000
$\mu\bar{\mu}$ Drell–Yan + $t\bar{t}$ $\sim$1,325,000 $\sim$1,260,000 $\sim$65,000
$\tau\bar{\tau}$ Training Validation Testing
@@
$\mu\bar{\mu}$ Training Validation Testing
Sideband 882,000 378,000 -
Signal Region
27,500 Simulated
82,500 Synthetic
27,500 Simulated
82,500 Synthetic
10,000 Background
10,000 Signal
@@ A.2 Generative Model
Generative models with conditional flow matching (CFM) use conditional normalizing flows by learning vector fields between conditional distributions. CFM avoids the computational intractability of traditional flow matching objectives [85] while offering improved efficiency and stability over diffusion models [90, 85].
Our flow is parameterized by a multi-layer perceptron with 6 hidden layers of 128 neurons each, using ELU activation functions [91]. Time is encoded using sinusoidal embeddings with 3 frequencies, producing a 6-dimensional time representation through concatenated sine and cosine components. The network takes as an input the concatenation of the features $\vec{x}$, time embeddings $t$ and conditioning variable $m$ (di-$\tau$ or di-$\mu$ visible invariant mass).
Training employs the Adam optimizer [92] with a cosine annealing learning rate scheduler, starting at $1\times 10^{-3}$ and decaying to $1\times 10^{-6}$. We used batch sizes of 1024 (di-$\mu$) and 512 (di-$\tau$). We trained the model over 2,000 to 5,000 epochs using A100 GPUs. We found that training beyond 5,000 epochs did not improve the quality of the generated samples. We fine tuned the hyperparameters of the generative model by minimizing the ROC between background and generated samples. Training for 5,000 epochs takes 32 hours.
After training, we select the model state with the lowest validation loss for the subsequent interpolation and sampling steps. The training and validation losses of the di-$\mu$ background can be seen in Fig. 6.
To generate synthetic background events in the signal region, we employ a two-step process that requires the trained flow to extrapolate beyond its training domain: (1) We fit a kernel density estimate to the di-$\ell$ invariant mass distribution observed in the signal region in the training set using a Gaussian kernel with bandwidth 0.01. (2) We sample mass values $m$ from this signal region mass distribution and use them to condition the learned normalizing flow that was trained on sideband data. This approach requires the learned normalizing flow to extrapolate from sideband masses (training domain) to signal region masses (target domain), assuming the conditional dependence on $m$ generalizes smoothly across this mass range.
@@
We do not model the hypersphere coordinates in this analysis due to their non-Euclidean nature, where normalizing flows are designed for features in the d-dimensional Euclidean space $\mathcal{R}^{d}$  [93]. The flow struggled to learn how to model the connection between different coordinates needed for a good background distribution. The simplex still captures much of the event’s topology, but the additional azimuthal information contained in the hypersphere could improve identification. There are limited works looking into flows on non-Euclidean manifolds [94, 95, 96]. However, implementing and improving upon their architectures for 10-dimensional hyperspheres is beyond the scope of this paper.
Figure 6: Training and validation loss for the di-$\mu$ channel, as an example.
Appendix B Comparison to $p_{T}$-based features
To evaluate the effectiveness of the phase space representation, we conducted a comparative analysis replacing the 10 simplex coordinates used with the $p_{T}$ values of the 10 highest $p_{T}$ objects in each event, in descending order. We focus on the di-$\mu$ channel as a proof of concept for this comparison.
To maintain consistency with the phase space analysis, events with fewer than 10 objects were zero-padded using randomized exponentially decreasing values from 0 to 25 GeV, corresponding to the minimum $p_{T}$ threshold in Delphes simulation. We retain the same complementary kinematic features used in the phase space analysis: the scalar sum of jet $p_{T}$, the missing transverse momentum (MET), and the angular separation $\Delta R$ between the two muons. These features provide essential event-level information not captured by the ranked $p_{T}$ values alone.
The features can be seen in Figure 7.
The same generative model architecture and training methodology employed for the phase space di-$\mu$ analysis were applied to the $p_{T}$-based features. The background generation achieved a ROC score of 0.53, comparable to the phase space background model performance (ROC = 0.52), indicating similar quality in modeling the background distribution.
Figure 8 shows the resulting significance improvement curves comparing the $p_{T}$ and phase space approaches. The phase space coordinates consistently outperform the $p_{T}$-based features across all signal injection levels.
The superior performance of phase space coordinates is encouraging for several reasons. First, the simplex coordinates are specifically designed to capture the topological structure of multi-particle final states in a coordinate-invariant manner, whereas $p_{T}$ rankings only preserve energy ordering without topological information. Second, our analysis utilizes only the simplex coordinates from the phase space decomposition and we have not yet incorporated the hypersphere coordinates, which encode additional angular correlations. This suggests significant untapped potential for further improvements in anomaly detection sensitivity using the complete phase space representation. The consistent performance advantage of phase space coordinates across different signal models demonstrates the robustness of this approach and supports its adoption for anomaly detection in collider physics applications.
Figure 7: A comparison of the distributions of true background events in the signal region and the synthetic ones generated by our normalizing flow for di-$\mu$. Object 10 $p_{T}$ is omitted for brevity, but has similar distributions.
Figure 8: A summary of the sensitivity enhancement of our anomaly detection approach for di-$\mu$ using phase space features (solid line) and $p_{T}$ features (dashed line). The x-axis shows the size of the signal, quantified as $\frac{S}{\sqrt{B}}$ prior to any anomaly detection selection. The y-axis shows the approximate signal significance achieved after the application of an anomaly score cut that is 1% efficient on the background. Our procedure utilizing phase space features is superior to using the $p_{T}$ of the objects alone in all signal cases.